### v1

# RuleR: Improving LLM Controllability by Rule-based Data Recycling

Abstract
Large language models (LLMs) still lack delicate controllability over their responses, which is critical to enhancing their performance and the user experience. However, curating supervised fine-tuning (SFT) datasets to improve LLM controllability usually relies on human experts or proprietary LLMs, which requires additional costs. To bridge this gap, we propose Rule-based Data Recycling (RuleR), a data augmentation method incorporating multiple constraints into the original data samples according to predefined rules, which creates new training tasks to consolidate the controllability of LLMs. Instead of creating new data from scratch, RuleR “recycles” existing data by simply applying rule-based edits to their responses and appending the rule-instructions in their original instructions. Experimental results demonstrate RuleR’s effectiveness in improving LLM controllability while maintaining general instruction-following capabilities. The code will be released on https://github.com/MingLiiii/RuleR.
RuleR: Improving LLM Controllability by Rule-based Data Recycling
Ming Li*1, Han Chenthanks: Equal Contribution., Chenguang Wang*2, Dang Nguyen1, Dianqi Li, Tianyi Zhou1 1University of Maryland 2Stony Brook University {minglii, tianyi}@umd.edu
1 Introduction
Figure 1: Comparing existing methods (top) and our RuleR (bottom) for enhancing LLM controllability. Most existing methods rely on extra-human/model supervision to generate or edit instructions and responses, neglecting the remaining potential of the original data. On the contrary, RuleR demonstrates that simple rule-based (human/model-free) editing of existing data can greatly improve LLM controllability.
Despite the remarkable advancement of the current LLMs and the continuous efforts to build high-quality supervised fine-tuning (SFT) datasets Brown et al. (2020); OpenAI (2023); Touvron et al. (2023a, b); Jiang et al. (2023), one critical aspect is how these LLMs better interact with human responses to maximize their utility and effectiveness for end-users Liu et al. (2024). According to the systematic investigation from Liu et al. (2024), it is essential for LLMs to constrain their outputs to follow specific user-defined formats or standards. Without constraints on responses, it is hard to directly put the free-formed responses into practical usage. This is also verified in the area of LLM Agents Li et al. (2024f); Chen et al. (2023, 2024b); Zhang et al. (2024), in which a predefined format is always necessary.
However, existing methods for building SFT datasets focus mainly on general instructions Wei et al. (2022); Wang et al. (2022); Taori et al. (2023); Xu et al. (2023); Zhou et al. (2023a); Li et al. (2023a); Zhang et al. (2023); Xu et al. (2024), neglecting the potential user-specific constraints and thus lacking delicate controllability of the response on user-desired lengths or formats Chen et al. (2024a); Xia et al. (2024). One potential method to enhance existing SFT data with multi-constraint controllability is to rewrite or modify instructions and responses by expert humans/models Xu et al. (2023); Li et al. (2023a, 2024b, 2024a); He et al. (2024) as shown in Figure 1 (top). However, these methods directly discard or rewrite the original data. Their data curation is costly and inefficient, requiring careful editing by human experts or proprietary LLMs. This raises the question: Can we “recycle” existing SFT datasets for controllability tuning by enforcing various types of constraints to them without human/model editing?
Drawing inspiration from recent work like Mosaic-IT Li et al. (2024c), which develops a human/model-free method augmenting SFT data, and IFEval Zhou et al. (2023b), which utilizes verifiable constraints to evaluate LLMs’ controllability, we propose Rule-based Data Recycling (RuleR), which automatically edits existing SFT data for improving LLM controllability. The key insight of RuleR is to replace human/model editing with various predefined rules of constraining patterns. Specifically, we manually inspect and construct several verifiable constraint rules, in which the original instructions are appended with various constraints, and the responses are optionally modified to fulfill these constraints based on their original characteristics. Our predefined rules cover a wide range of diverse constraints ranging from high-level constraints, e.g., controlling the word frequency in the response, to lower-level constraints, e.g., setting specific wrapping formats [^1] The complete list of rules and descriptions can be found in the Appendix C. . Each rule contains (1) a set of constraint templates that serve as additional rule-instructions, on constraints the LLM should follow, and (2) specified operations that alternately edit the instruction and response to reach an alignment between them (i.e., all the constraints in the instruction hold in the response). For each sample from the original dataset, we randomly draw several rules to be applied in the editing. This produces an augmented sample with the constraints enforced that can be used for controllability tuning.
An example is shown in Figure 1 (bottom). The original instruction is “What type of fruit would be a good addition to a fruit salad?”, and the corresponding response is “Grapes, kiwis, oranges, apples.”. If the constraint “Num of Nouns” is sampled and implemented on this sample, the augmented instruction becomes “What type of fruit would be a good addition to a fruit salad? Ensure the response has more than 3 nouns.”, and the original response remains unchanged on this type of rule. In this example, the constraint (number of nouns) is directly decided by the original data sample (4 nouns), thus avoiding the regeneration of response.
Extensive experimental results on the IFEval (Instruction-Following Eval) Zhou et al. (2023b) benchmark on various datasets demonstrate the effectiveness of our human/model-free method in enhancing LLM controllability. Moreover, the instruction-following metrics also show that RuleR still preserves the general instruction-following ability promoted by the original SFT dataset. To the best of our knowledge, RuleR is the first human/model-free data augmentation approach designed to improve LLM controllability in enforcing multiple constraints to LLM-generated response.
2 Method
2.1 Preliminaries
Given a supervised finetuning dataset $D$, there are $n$ data samples each of which is represented by a tuple $(x_{i},y_{i})$, where $x_{i}$ represents the instruction and $y_{i}$ represents the corresponding response. Let $p_{\theta}(\cdot)$ denote the LLMs to be trained, with parameters $\theta$. In the instruction tuning setting, $p_{\theta}$ is typically fine-tuned by maximizing the following objective on each data $(x_{i},y_{i})$, $y_{i,j}$ represents the $j_{th}$ token of response $y_{i}$ and $l_{i}$ represents the token length:
$\max_{\theta}\sum_{j=1}^{l_{i}}\log p_{\theta}\left(y_{i,j}|x_{i}\right),$ (1)
2.2 Rule-based Data Recycling
The basis for augmenting existing SFT datasets is to reconstruct instructions and responses. Most existing methods utilize human experts or strong teacher models for this process, which is costly, inefficient, and fails to get fully used to the existing data, as represented in Figure 1 (top). Inspired by Mosaic-IT which proposes a human/model-free augmentation method to reconstruct instruction-response pairs, we are motivated to design a rule-based method for controllability augmentation.
Model Dataset Method Instruction Following Eval Pair-wise Open LLM Leaderboard
Prompt (S) Inst (S) Prompt (L) Inst (L) Winning Score Average A H M T
Mistral-7B Alpaca-GPT4 Baseline 32.72 42.45 35.67 45.44 1.000 61.24 56.23 81.07 56.22 51.42
Ours 39.56 51.44 43.44 55.40 1.044 61.05 56.66 80.50 57.24 49.81
Alpaca Baseline 33.64 44.60 36.23 47.96 1.000 55.15 51.96 74.61 52.85 41.20
Ours 35.12 46.76 37.89 49.52 1.158 56.21 54.61 77.70 54.54 38.00
Wizard-70k Baseline 37.34 48.68 40.85 52.04 1.000 59.38 54.61 79.96 55.68 47.27
Ours 46.77 57.07 49.17 59.71 1.168 59.75 55.38 80.75 55.59 47.27
Recycled Wizard Baseline 30.87 42.41 35.86 46.40 1.000 59.61 54.10 77.85 57.61 48.87
Ours 39.56 51.08 45.10 56.24 0.987 60.43 56.66 78.01 58.61 48.43
DEITA 6K Baseline 41.22 51.08 44.55 54.92 1.000 64.82 60.41 82.52 61.57 54.76
Ours 45.10 54.32 47.69 56.95 1.010 65.34 61.16 82.78 62.74 54.66
Llama2-7B Alpaca-GPT4 Baseline 26.25 36.33 30.31 40.29 1.000 58.71 54.69 80.05 47.89 52.21
Ours 32.35 42.09 35.30 45.56 1.070 59.77 56.74 80.67 48.45 53.21
Alpaca Baseline 31.42 40.77 33.46 43.17 1.000 55.25 54.35 78.65 47.02 40.98
Ours 34.38 44.36 37.34 47.36 1.023 55.24 54.61 78.76 46.17 41.42
Wizard-70k Baseline 31.24 44.24 35.49 48.68 1.000 57.09 54.18 79.25 46.93 48.02
Ours 38.82 50.12 42.33 53.48 1.087 57.25 55.20 79.81 46.61 47.38
Llama2-13B Alpaca-GPT4 Baseline 32.90 44.60 36.23 48.08 1.000 61.47 58.70 83.12 54.13 49.92
Ours 36.60 47.00 37.89 49.28 0.977 61.96 59.47 82.88 53.98 51.52
Alpaca Baseline 34.94 44.36 36.41 46.29 1.000 57.63 57.25 81.23 54.13 37.91
Ours 36.04 48.20 41.22 52.88 0.977 57.16 57.17 81.11 52.70 37.65
Wizard-70k Baseline 43.07 53.84 46.40 57.67 1.000 61.24 57.04 83.39 55.76 48.78
Ours 45.47 58.15 50.09 61.99 1.010 60.84 58.28 82.37 54.35 48.36
Table 1: The comparisons on the Instruction Following Eval, Pair-wise Comparison Winning Score, and the Open LLM Leaderboard, on 3 different base models and several different instruction tuning datasets. The baselines are models trained with the original dataset. Prompt and Inst represent Prompt-level and Instruction-level accuracy; S and L represent Strict and Loose versions. A, H, M, and T represent ARC, HellaSwag, MMLU, and TruthfulQA.
With this goal, the first challenge is the reformulation of instructions. Due to the lack of human experts or LLMs to generate the corresponding responses for any given instructions, it is not feasible to add arbitrary constraints to the original instructions. To solve this problem, we propose to only incorporate constraints that are compatible with the original data sample. Specifically, we propose to enhance LLM controllability by Rule-based Data Recycling (RuleR), in which the rules and constraints used to reformulate instruction are determined by the original data, ensuring the alignment between instruction and response. Motivated by IFEval, we manually inspect and propose several verifiable rules in which the responses are not required for regeneration by LLMs or human experts, but only need optional slight adjustments. The complete list of rules is listed in Table 3 ranging from high-level constraints like term frequency to lower-level constraints like specific wrapping formats.
During implementation, we instantiate each predefined rule as a tuple , $(\mathbf{S}_{k},f_{k},g_{k})$, where $\mathbf{S}_{k}$ represents the set of manually curated rule-instruction templates for the $k^{th}$ rule, $f_{k}$ and $g_{k}$ represent the function to reformulate instructions and optionally modify responses respectively, as shown in Figure 1(bottom). For each data sample $(x_{i},y_{i})$, the augmented instruction $x_{i,aug}$ will be reformulated according to the characteristics of the original data and the corresponding template sets:
$x_{i,aug}=f_{k}(x_{i},y_{i},\mathbf{S}_{k})$
where specifically, a random rule-instruction template will be sampled and filled with the corresponding characteristics as the additional constraint to the original instruction. Then the rule-instruction is concatenated with the original instruction becoming the augmented instruction. Similarly, the augmented response $y_{i,aug}$ will be optionally modified based on the selected rule $k$:
$y_{i,aug}=g_{k}(x_{i},y_{i},\mathbf{S}_{k}),$
where the function $g_{k}$ varies dramatically for different rules: For some of the rules, no modification needs to be done to the original response, while for others, some arbitrary part of the response needs to be replaced with the desired user format.
3 Experiments
3.1 Experimental Setup
We conduct experiments on several datasets including the Alpaca Taori et al. (2023), Alpaca-GPT4 Peng et al. (2023), WizardLM Xu et al. (2023), Recycled WizardLM Li et al. (2023a), DEITA Liu et al. (2023), and on several base LLMs including Llama2-7B, Llama2-13B Touvron et al. (2023b), and Mistral-7B Jiang et al. (2023), introduction on datasets can be found in the Appendix A. We utilize the prompt and code base from Vicuna Chiang et al. (2023a) and flash attention Dao et al. (2022). The Adam optimizer Kingma and Ba (2017) is utilized with the batch size of $128$ and with the max token length of $2048$. For training on Llama2-7B and Llama2-13B Touvron et al. (2023b), the maximum learning rate is set to $2\times 10^{-5}$ with a warmup rate of $0.03$ for $3$ epochs. For training on Mistral-7B Jiang et al. (2023), the maximum learning rate is set to $1\times 10^{-5}$ with a warmup rate of $0.1$ for $2$ epochs. When utilizing our method, we run the augmentation process $3$/$2$ times to simulate the epochs of training. These augmented data are then mixed together and used for training $1$ epoch. All other configurations are kept the same as the baselines.
3.2 Experimental Results
The main experimental results are presented in Table 1, containing the performance comparison on the Instruction Following Eval Zhou et al. (2023b), Pair-wise Comparison Winning Score, and the Open LLM Leaderboard Gao et al. (2021a), on 3 different base models and several different instruction tuning datasets. The Pair-wise Comparison Winning Score is calculated as (Num(Win)--Num(Lose))///Num(All) $+1$ and the values that are greater than $1.0$ represent better responses generated. Detailed descriptions of evaluation metrics can be found in the Appendix B.
Compared to the baseline, our method has consistent improvements on the IF Eval benchmark, across different base models and datasets, which aims at measuring LLMs’ constraint-following abilities by using verifiable instructions. It is astonishing that our method can improve the IF Eval scores by approximately $10\%$ on some of the configurations, by just utilizing the rule-based recycling method on the original data, without any human/model edition. Moreover, the performances keep being positive on originally diverse and high-quality datasets like Recycled WizardLM Li et al. (2023a) and DEITA Liu et al. (2023), which further verify the potential of our method. Compared with existing methods which enhance LLMs’ constraint controllability by generating totally new data, our method focuses more on fully utilizing the potential of existing data.
Furthermore, our method not only improves the constraint controllability but also keeps the general instruction-following ability of the original data. The Pair-wise comparison and Open LLM leaderboard results showcase comparable or sometimes better performances compared with the baseline models. We hypothesize that the additional constraints might complicate the original instructions, thus leading to potentially improved instruction-following abilities.
3.3 Ablation Studies
Evaluation Metrics IF Eval Pair-wise Open LLM
Baseline 39.07 1.000 61.24
Single Temp. 42.62 0.987 61.43
Diverse Temp. (*) 47.46 1.044 61.05
Max Rule = 1 46.14 1.168 61.22
Max Rule = 2 47.09 1.117 61.15
Max Rule = 3 (*) 47.46 1.044 61.05
Max Rule = 4 46.89 1.003 61.55
Max Rule = 5 44.36 1.013 60.06
Aug Rate = 0.1 41.72 1.020 61.34
Aug Rate = 0.3 42.08 1.037 61.20
Aug Rate = 0.5 46.55 1.111 61.31
Aug Rate = 0.7 46.23 1.111 61.18
Aug Rate = 0.9 (*) 47.46 1.044 61.05
Table 2: Ablations. The experiments are conducted on Mistral-7B on Alpaca-GPT4 data and the configuration notated with “(*)” represents our default setting. “IF Eval” and “Open LLM” represent the averaged scores.
In this section, ablation experiments are conducted on Mistral-7B with the Alpaca-GPT4 dataset.
As shown in Table 2, “Single Temp.” represents utilizing only one rule-insturtcion template for each rule, while “Diverse Temp.” represents utilizing several different templates (with the same meaning) for each rule and those templates are randomly sampled during the augmentation. The superiority of “Diverse Temp.” showcases the necessity of providing diverse templates to avoid potential overfitting. “Max Rule = x” represents the setting in which at most x different rules can be sampled and utilized on each original data sample. The performances of models, whose Max Rule is set under 5, are promising, while the performance declines on all 3 evaluation metrics when the Max Rule is set as 5. These results indicate that too many rules utilized together in one sample might be harmful to LLM capability. “Aug Rate = x” represents there is a probability of x to apply our augmentation to each sample. It can be observed that increasing the augmentation rate is beneficial to the controllability of LLMs, while slightly harmful to the general instruction-following capability. However, the gaps on IF Eval are much larger than the other 2 metrics, indicating this rate will mostly influence the multi-constraint controllability of LLM.
4 Conclusion
In this work, we proposed a novel method, Rule-based Data Recycling (RuleR) which utilizes the constraints derived from the existing data to enhance the dataset and incorporates these constraints into the training process, thereby improving the multi-constraint controllability of LLMs. This approach presents a promising direction for developing SFT-augmenting methods to improve LLM controllability, paving the way for more efficient advancements of LLM.
Limitations
Our method focuses on improving LLM controllability by rule-based data augmentation to avoid extra regeneration processes from humans or models. Though the alleviation of reliance on human/model editing is our main contribution, it inevitably limits the types of constraints that can be implemented on the original data sample. Currently, all the constraint rules are based on verifiable shallow syntactic characteristics like the occurrence and frequency of words or sentences, etc, lacking constraint rules that focus more on the semantic feature or content. Thus, our method shows great potential to be further enhanced if further involves small models, which will still remain efficient and avoid the regeneration process.
Appendix A Training Dataset
We utilize 5 SFT datasets to evaluate the effectiveness of our method:
Alpaca dataset Taori et al. (2023): This dataset consists of 52,000 instruction-following samples created using the self-instruct paradigm Wang et al. (2023b) and OpenAI’s text-davinci-003 model. Characterized as a classical dataset with moderate-quality attributes, it serves as the fundamental validation.
Alpaca-GPT4 dataset Peng et al. (2023): This dataset is an enhanced Alpaca dataset that includes responses generated by GPT-4.
WizardLM dataset Xu et al. (2023): This dataset is generated by the novel Evol-Instruct method, which utilizes ChatGPT-3.5 to rewrite instructions step by step into more complex ones and generate the corresponding responses. We utilize the 70k version in our method, which comprises 70,000 high-quality SFT samples.
Recycled WizardLM Dataset Li et al. (2023a): This dataset is an improved version of the WizardLM dataset, by utilizing the Reflection-Tuning method. In the Reflection-Tuning, the initial dataset undergoes two main phases: Reflection on Instruction and Reflection on Response. In the first phase, specific criteria are carefully curated to evaluate and refine the initial instructions. During the second phase, responses are thoroughly examined and improved to align with the refined instructions. This process generates a dataset with superior quality compared to the original dataset.
DEITA dataset Liu et al. (2023): This dataset leverages the DEITA (Data-Efficient Instruction Tuning for Alignment) method to select high-quality data from a pool comprised of several high-quality datasets, such as WizardLM and Alpaca. DEITA employs a score-first, diversity-aware data selection strategy to optimize the selection process. This strategy uses a GPT-as-a-judge scoring system that combines complexity and quality in a practical and straightforward manner. The scores are incorporated with the diversity-based selection, ensuring that all the data maintains high standards of complexity, quality, and diversity.
Appendix B Evaluation Metrics
We employ 3 commonly accepted metrics for the evaluation, including IFEval (Instruction-Following Eval), Pair-wise Comparison, and Open LLM Leaderboard
IFEval Zhou et al. (2023b) is the primary evaluation metric employed in our study due to its compatibility with our motivation. It focuses on evaluating how LLMs follow various additional constraints, such as specifying a word count or requiring the inclusion of certain keywords a specified number of times. To avoid the utilization of LLMs during evaluation, it proposes 25 distinct types of verifiable instructions. There are 541 prompts in total and each of them incorporates one or more of these verifiable instructions, ensuring the comprehensiveness of the evaluation. IFEval serves as a great inspiration for our method, and there exist semantical overlappings between their verifiable instructions and our rules. However, the specific prompts used in IFEval are kept unknown in the construction of our rule templates, avoiding potential data leakage. Moreover, IFEval only needs to verify the responses, while our method needs to modify the responses for the training, which pushes this process a step further. Consequently, this benchmark not only facilitates a comprehensive comparison but also provides valuable insights that align with our purpose.
Pair-wise Comparison involves evaluating responses from LLMs like GPT-4, especially in open-domain contexts. This method has shown a notable alignment with human assessments, providing a credible evaluative foundation Zheng et al. (2024); Li et al. (2023b); Sottana et al. (2023). We utilize test instruction sets from WizardLM Xu et al. (2023) and Vicuna Chiang et al. (2023b), comprising 218 and 80 diverse human-curated instructions, respectively. Following the framework by Li et al. (2024e, d), we prompt judging LLM to rate responses on a scale from 1 to 10 across multiple dimensions such as relevance and accuracy. To mitigate positional bias Ko et al. (2020); Wang et al. (2023a), responses are presented in two distinct sequences. Evaluation outcomes are categorized into the ’win-tie-lose’ system for each instruction. One model’s response can be noted as Win/Lose only if it has a Higher/Lower score in both directions. The specific evaluation prompt is presented in Figure 2.
Open LLM Leaderboard employs Eval Harness Gao et al. (2021a) to systematically assess generative language models through diverse tasks. Key benchmarks include ARC Clark et al. (2018), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2020), and TruthfulQA Lin et al. (2021), collectively evaluating reasoning abilities, common-sense knowledge, and factual accuracy. This approach highlights LLM strengths and areas for improvement in critical aspects of language understanding and generation Gao et al. (2021b).
  Prompt for Performance Evaluation
  System Prompt
You are a helpful and precise assistant for checking the quality of the answer.
User Prompt
[Question]
Question
[The Start of Assistant 2’s Answer]
Answer 2
[The End of Assistant 2’s Answer]
[The Start of Assistant 2’s Answer]
Answer 2
[The End of Assistant 2’s Answer]
We would like to request your feedback on the performance of two AI assistants in response to the user question displayed above.
Please rate the helpfulness, relevance, accuracy, level of details of their responses. Each assistant receives an overall score on a scale of 1 to 10, where a higher score indicates better overall performance.
Please first output a single line containing only two values indicating the scores for Assistant 1 and 2, respectively. The two scores are separated by a space. In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgment.
Figure 2: The prompt we used to request GPT4 to evaluate the responses.
Appendix C Predefined Rules
Rule Type Rule Name Example Rule-Instruction Templates
Keyword Frequency Keyword Appearance Ensure {Keyword} is in the response.
Keyword Frequency Keyword Frequency Ensure there are more/less/exact {N} {Keyword} in the response.
Number Constraint Num of Adjectives Ensure the response has more/less/exact {N} adjectives.
Number Constraint Num of Nouns Ensure the response has more/less/exact {N} nouns.
Number Constraint Num of Verbs Ensure the response has more/less/exact {N} verbs.
Number Constraint Num of Characters Ensure the response has more/less/exact {N} characters.
Number Constraint Num of Letters Ensure the response has more/less/exact {N} letters.
Number Constraint Num of Words Ensure the response has more/less/exact {N} words.
Number Constraint Num of Sentences Ensure the response has more/less/exact {N} sentences.
Number Constraint Num of Paragraphs Ensure the response has more/less/exact {N} paragraphs.
Number Constraint Num of Bullets Ensure the response has more/less/exact {N} bullet points.
Repetition Instruction Repetition Repeat the instruction before providing the response.
Repetition Response Repetition Repeat the response {N} times.
Case All Up Case Ensure the response is all in upper case.
Case All Low Case Ensure the response is all in lowercase.
Case Target Letter Case Ensure all the letters {x} in the response are in uppercase.
Case Target Keyword Case Ensure all the word {Keyword} in the response are in uppercase.
Case Target Sentence Case Ensure {i}-th sentence in the response is in uppercase.
Case Target Paragraph Case Ensure {i}-th paragraph in the response is in uppercase.
Punctuation All All Removal Ignore all punctuation in the response.
Punctuation All All Replacement Use {Symbol} to replace all punctuation in the response.
Punctuation Target Target Removal Ignore {Punctuation} punctuation in the response.
Punctuation Target Target Replacement Use {Symbol} to replace {Punctuation} in the response.
Format Wrapping Keyword Wrapping Ensure every {Keyword} is wrapped in {Format} in the response.
Format Wrapping Sentence Wrapping Ensure {i}-th sentence is wrapped in {Format} in the response.
Format Wrapping Bullet Wrapping Ensure {i}-th bullet point is wrapped in {Format} in the response.
Format Wrapping Paragraph Wrapping Ensure {i}-th paragraph is wrapped in {Format} in the response.
Formatted Repeating Instruction Wrapping Repeat the instruction in {Format} before providing the response.
Formatted Repeating Response Wrapping Repeat the response {N} times in {Format}.
Table 3: The list of predefined constraint rules.
In this section, we will dive into the predefined rules describing each constraint specifically.
Keyword Appearance simulates the scenario where specific keywords are required to appear in the responses. In this rule, several non-stop words are randomly selected from the original data sample and used as the desired characteristics. The placeholder Keyword in the constraint template will be replaced by the sampled keyword as the rule-instruction. This process can be repeated to simulate the constraints on multiple keywords. The augmented instruction is the concatenation of the original instruction and rule-instruction. The original response does not need to be modified in this rule and is directly used as the augmented response.
Keyword Frequency simulates controlling the frequency of specific keywords in generated responses. In this rule, several non-stop words and their frequencies are randomly sampled and used as the desired characteristics. There are three random sub-situations in the rule: More, Less, or Equal. In the “Equal” situation, the placeholders {N} and {Keyword} will be directly replaced by the sampled keyword and its frequency. In the “More” or “Less” situations, a small random number $x$ will be randomly generated to adjust the keyword frequency to meet the desired constraint template, such as “Ensure there are more than {N - x} {Keyword}” or “Ensure there are fewer than {N + x} {Keyword}.” This process can be repeated to simulate constraints on multiple keywords. The augmented instruction is the concatenation of the original instruction and rule-instruction. The original response does not need to be modified in this rule and is directly used as the augmented response.
Num of Adjectives simulates controlling the total number of adjectives in generated responses. In this rule, the adjectives in the original response are identified and counted using part-of-speech tagging (POS). There are three possible sub-situations in the rule: More, Less, or Exact. In the “Exact” situation, the placeholder {N} will be replaced by the number of adjectives. In the “More” or “Less” situations, a small random number $x$ is randomly generated to adjust N to meet the constraints, such as “Ensure the response has more than {N - x} adjectives” or “Ensure the response has fewer than {N + x} adjectives.” This process can only be used once for each sample. The augmented instruction is the concatenation of the original instruction and rule-instruction. The original response does not need to be modified in this rule and is directly used as the augmented response.
Num of Nouns simulates controlling the number of nouns in generated responses, similar to the “Num of Adjectives”.
Num of Verbs simulates controlling the number of verbs in generated responses.
Num of Characters simulates controlling the number of characters in generated responses.
Num of Letters simulates controlling the number of letters in generated responses.
Num of Words This rule simulates controlling the number of words in generated responses.
Num of Sentences simulates controlling the number of sentences in generated responses. The sentences from the original response are segmented by utilizing dependency parsing.
Num of Paragraphs simulates controlling the number of paragraphs in generated responses. The paragraphs from the original response are segmented by regular expressions.
Num of Bullets simulates controlling the number of bullet points in generated responses. The bullet points from the original response are segmented by regular expressions.
Instruction Repetition simulates the scenario where the LLM is requested to repeat the instructions before providing the response. This process can be applied only once for each instruction. The augmented instruction is the concatenation of the original instruction and rule-instruction. The augmented response is the concatenation of repeated original instruction and the response.
Response Repetition simulates the scenario where the LLM is requested to repeat the responses several times. In this rule, the response is repeated {N} times, where {N} is a random number. This process can only be applied once per data sample. The augmented instruction is the concatenation of the original instruction and rule-instruction. The augmented response is the concatenation of N identical responses.
UP Case simulates requesting the entire response is required in uppercase. In this rule, the original response is converted to uppercase format entirely. This process can only be used once for each response. The augmented instruction is the concatenation of the original instruction and rule-instruction. The augmented response is the all-uppercase version of the original response.
Low Case simulates requesting the entire response is required in lowercase.
Letter Case simulates the scenario where specific types of letters in the response are required to be in uppercase. In this rule, the specific letter x is sampled from the response, and all occurrences of this letter in the response are capitalized. This process can be repeated on different random letters. The augmented instruction is the concatenation of the original instruction and rule-instruction. The augmented response is the original response with all specific letters in uppercase.
Keyword Case simulates the scenario where specific keywords in the response are required to be in uppercase. The specific keyword Keyword is sampled from the response.
Sentence Case simulates the scenario where the specific sentences in the response are required to be in uppercase. The index of the sentence i is randomly selected within the total number of sentences in the response.
Paragraph Case simulates the scenario where the specific paragraphs in the response are required to be in uppercase. The index of the paragraph i is randomly selected within the total number of paragraphs in the response
All Removal simulates controlling LLM to ignore the use of punctuation. In this rule, the punctuation marks in the response will be removed completely, and the new response will serve as the augmented response. This process can only be used once for each sample.
All Replacement simulates controlling LLM to replace all the original punctuation with a predefined symbol {Symbol}. In this rule, the punctuation marks in the response will be replaced, and the new response will serve as the augmented response. This process can only be used once for each sample.
Target Removal simulates the scenario where a specific type of punctuation mark {Punctuation} in the response is ignored. In this rule, a random type of punctuation is identified, and all occurrences of this mark are removed, the new response will serve as the augmented response. This process can only be used once for each sample to avoid confusion.
Target Replacement simulates the scenario where a specific type of punctuation mark {Punctuation} in the response is replaced by a specified symbol {Symbol}. In this rule, a random type of punctuation mark is identified, and all occurrences of this mark will be replaced by a predefined symbol in the response. This process can only be used once.
Keyword Wrapping simulates the scenario where specific keywords in the response are required to be wrapped in a specified format. In this rule, a randomly chosen {Keyword} is identified, and all occurrences of this keyword are wrapped in the randomly specified {Format} in the response. This process can be repeated several times on different words with various formats. The augmented instruction is the concatenation of the original instruction and rule-instruction. The augmented response is the original response with all keywords wrapped in the format.
Sentence Wrapping simulates the scenario where specific sentences in the response are required to be wrapped in a specified format. The index of the sentence i is randomly selected within the total number of sentences in the response.
Bullet Wrapping simulates the scenario where specific bullet points in the response are required to be wrapped in a specified format. The index of the bullet point i is randomly selected within the total number of bullet points in the response.
Paragraph Wrapping simulates the scenario where a specific paragraph in the response are required to be wrapped in a specified format. The index of the paragraph i is randomly selected within the total number of paragraphs in the response.
Instruction Wrapping simulates a scenario where the original instruction is required to be repeated in a specified format before providing the response. In this rule, the original instruction is restated with wrapping in the randomly chosen {Format} before giving the actual response. This process can be applied only once for each instruction. The augmented instruction is the concatenation of the original instruction and rule-instruction. The augmented response is the concatenation of the original instruction wrapped in the specific format and the response.
Response Wrapping simulates a scenario where the response wrapped in the specific format is required to be repeated several times. In this rule, the response is repeated {N} times wrapeed in the specified {Format}, with {N} and {Format} being randomly selected. This process can be applied only once. The augmented instruction is the concatenation of the original instruction and rule-instruction. The augmented response is the concatenation of the repeated response wrapped in the format.